# Wavelet Enhanced Adaptive Frequency Filter for Sequential Recommendation

### v1

@@ Abstract
Sequential recommendation has garnered significant attention for its ability to capture dynamic preferences by mining users’ historical interaction data. Given that users’ complex and intertwined periodic preferences are difficult to disentangle in the time domain, recent research is exploring frequency domain analysis to identify these hidden patterns. However, current frequency-domain-based methods suffer from two key limitations: (i) They primarily employ static filters with fixed characteristics, overlooking the personalized nature of behavioral patterns; (ii) While the global discrete Fourier transform excels at modeling long-range dependencies, it can blur non-stationary signals and short-term fluctuations. To overcome these limitations, we propose a novel method called Wavelet Enhanced Adaptive Frequency Filter for Sequential Recommendation (WEARec). Specifically, it consists of two vital modules: dynamic frequency-domain filtering and wavelet feature enhancement. The former is used to dynamically adjust filtering operations based on behavioral sequences to extract personalized global information, and the latter integrates wavelet transform to reconstruct sequences, enhancing blurred non-stationary signals and short-term fluctuations. Finally, these two modules work synergistically to achieve comprehensive performance and efficiency optimization in long sequential recommendation scenarios. Extensive experiments on four widely-used benchmark datasets demonstrate the superiority of WEARec. Our code is available at https://github.com/xhy963319431/WEARec.
@@ Introduction
Sequential Recommendation (SR) plays a crucial role in e-commerce applications by capturing users’ dynamic interest shifts through their historical interaction data (Schedl et al. 2018; Hansen et al. 2020). The remarkable success of the transformer architecture in Natural Language Processing (NLP) (Vaswani et al. 2017) and Computer Vision (CV) (Carion et al. 2020; Dosovitskiy et al. 2020) has led to significant advancements in sequential recommendation (Kang and McAuley 2018; Xu et al. 2021; Zhang et al. 2019; Liu et al. 2021). This has directly inspired a multitude of sequential recommendation models based on self-attention (Sun et al. 2019; Qiu et al. 2022; Zhou et al. 2020). However, items in user interactions are typically chronologically entangled and inherently noisy (Du et al. 2023a, b). Consequently, it is challenging for models to directly discern changes in behavioral preferences from raw sequences within the temporal domain (Zhou et al. 2022; Xu et al. 2019).
@@
To address this limitation, recent research has begun exploring frequency-domain approaches to replace self-attention mechanisms (Lee-Thorp et al. 2021; Rao et al. 2021; Fein-Ashley, Kannan, and Prasanna 2025; Zhou et al. 2024). By decomposing user sequences into different frequency components ($e.g.$, high-frequency and low-frequency signals) using Fourier transform, periodic patterns that are difficult to identify in the time domain can be effectively captured (Zhou et al. 2022; Du et al. 2023a; Shin et al. 2024; Zhang et al. 2025). For example, FMLPRec (Zhou et al. 2022) pioneered frequency domain processing of sequential data, replacing the self-attention mechanism with learnable filters. SLIME4Rec (Du et al. 2023a) proposed a frequency ramp structure, which considers different frequency bands for each layer. BSARec (Shin et al. 2024) used a frequency domain re-tuning component as an inductive bias for self-attention.
@@
However, despite their success in SR, existing frequency-domain sequential recommendation models have two key limitations. First, existing methods typically apply a static, fixed-pattern filter to all frequency components, which uniformly processes all user sequences, ignoring the personalized nature of behavioral patterns (Shin et al. 2024; Zhang et al. 2025). This unified filtering approach is susceptible to the influence of dominant users in the dataset and fails to consider the characteristics of individual users. In fact, users often exhibit diverse behavioral patterns: some user behaviors follow long-term preferences ($e.g.$, low-frequency signals), while other user behaviors show the opposite trend (Du et al. 2023a; Zhang et al. 2025). To illustrate this, we trained a classic sequential recommendation model ($i.e.$, FMLPRec (Zhou et al. 2022)) on the Beauty and Sports datasets, and replaced its learnable filters with specific band-pass filters. By statistically analyzing how many users could be correctly predicted by specific frequency components, we could identify which users were driven by particular frequency components. The results, shown in Figure 1, indicate that users exhibit diverse behavioral patterns, with each focusing on different frequencies. This emphasizes the importance of developing personalized filtering models to capture individualized user behavioral patterns.
@@
The second limitation is related to the low-pass filtering characteristic of frequency-domain filters. The Discrete Fourier Transform (DFT) analyzes signal components globally, primarily serving as a global, rather than local, frequency extraction method. While the global DFT excels at capturing long-range dependencies in current frequency-domain recommendation models (Zhou et al. 2022; Du et al. 2023a), it struggles to capture the local temporal features of high-frequency interactions and short-term points of interest (Lu et al. 2025). For instance, FMLPRec has been shown to essentially act as a low-pass filter (Shin et al. 2024). Although SLIME4Rec attempts to balance high-frequency and low-frequency representations through its hierarchical learning mechanism, the model still tends to learn low-frequency components within the hierarchical frequency bands.
@@
To overcome these challenges, we propose Wavelet Enhanced Adaptive Frequency Filter for Sequential Recommendation (WEARec). WEARec consists of two key modules: Dynamic Frequency Filtering (DFF) and Wavelet Feature Enhancement (WFE) module. Specifically, the DFF module uses a simple Multi-Layer Perceptron (MLP) to enhance or suppress specific frequency bands based on context signals, ensuring effective global fusion. Furthermore, the WFE module reconstructs sequences via wavelet transform, amplifying obscure non-stationary signals and short-term fluctuations that are prone to being blurred by global DFT. Finally, to ensure all frequency components are considered and to best preserve meaningful periodic user features, we blend the DFF module with the WFE module. Moreover, our proposed model achieves better performance with lower computational costs, especially in long-sequence scenarios. The main contributions of this paper are summarized as follows:
@@
Where $\tilde{L}$ and $\tilde{H}$ are reconstruction filters. In our paper, the forward DWT converts sequential behavior into high/low-frequency sub-signals and denote it by $\mathcal{W}(\textperiodcentered)$. The IDWT (denoted by $\mathcal{W}^{-1}(\textperiodcentered)$) reconstructs decomposed sub-signals into the original signal. For more descriptions, interested readers should refer to Appendix A.
@@
This module captures fine-grained temporal patterns through differentiable wavelet transforms. Here, the Haar wavelet transform (Stanković and Falkowski 2003) was selected due to its simple structure, high computational efficiency, and the desirable property of perfect signal reconstruction.
@@
In the final layer of WEARec, we can compute the recommendation probability for each candidate item to predict how likely the user would adopt the item. Specifically, the corresponding predicted probability $\mathbb{\hat{y}}$ can be generated by:
@@
where $\hat{\mathbb{y}}\in\mathbb{R}^{|\mathcal{V}|}$. and $\mathbb{h}^{L}\in\mathbb{R}^{1\times d}$ is the output of the $L$-layer blocks at the final step. To optimize the model parameters, we therefore use the cross-entropy loss (Qiu et al. 2022; Du et al. 2023b, a; Shin et al. 2024). The objective function of SR can be formulated as:
@@
Datasets. We conduct experiments on four public datasets collected from real-world platforms in order to thoroughly evaluate WEARec. i,ii) Beauty and Sports from Amazon (McAuley et al. 2015), iii) ML-1M (Harper and Konstan 2015), iv) LastFM. Following (Zhou et al. 2020, 2022) , we also adopt the 5-core settings by filtering out users with less than 5 interactions. The detailed dataset statistics are presented in Appendix B.1.
Evaluation Metrics. In our evaluation, we adopt the leave-one-out strategy for partitioning each user’s item sequence (Zhou et al. 2020). We rank the prediction scores throughout the entire item set without using negative sampling, as recommended by (Krichene and Rendle 2020). Performance is evaluated on a variety of evaluation metrics, including Hit Ratio at $K$ (HR@$K$) and Normalized Discounted Cumulative Gain at $K$ (NDCG@$K$, NG@$K$) on all datasets. The $K$ is set to 10 and 20.
@@
Time-domain SR models: GRU4Rec (Hidasi et al. 2015), Caser (Tang and Wang 2018), SASRec (Kang and McAuley 2018), and DuoRec (Qiu et al. 2022).
Frequency-domain SR models: FMLPRec (Zhou et al. 2022), FamouSRec (Zhang et al. 2025), FEARec (Du et al. 2023b), SLIME4Rec (Du et al. 2023a), BSARec (Shin et al. 2024).
Implementation Details. We implement our WEARec model in PyTorch. All the experiments are conducted on a Linux server equipped with a 32GB NVIDIA Tesla V100 GPU. For the baseline models, we refer to their best hyper-parameters setups reported in the original papers and directly report their reimplementations results if available, since the datasets and evaluation metrics used in these works are strictly consistent with ours. Both the dimension of the feed-forward network and item embedding size are set to 64. The number of WEARec blocks $L$ is set to 2, and the maximum sequence length $N$ is set to 50. Batch size is set to 256. The model is optimized by Adam optimizer with a learning rate from {0.0005,0.001}. The wavelet decomposition level is set to 1. The $\alpha$ is in {0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9}, and the $k$ chosen from {1,2,4,8}. We report the result of each model under its optimal hyper-parameter settings. The best hyperparameters are in Appendix B.3 for reproducibility.
@@
Table 2: The number of parameters and training time (runtime per epoch) for $N=200$ on ML-1M and LastFM. More results in other datasets are in Appendix B.
Model Performance. To investigate the impact of long sequence scenarios on recommendation results, we varied the maximum sequence length $N$ for FMLPRec, BSARec, SLIME4Rec and WEARec. We selected the LastFM and ML-1M datasets, which have longer average sequence lengths, for our experiments. Figure 3 presents the experimental results in terms of HR@20. We have obtained similar experimental results in terms of other metrics. We observed that almost all models achieved their best performance at $N=200$, indicating that longer sequence information can more comprehensively represent user behavior patterns. Furthermore, while baseline models showed performance improvements in long-sequence scenarios, they were prone to overfitting, leading to performance convergence. Finally, WEARec consistently outperformed the baselines across all different maximum sequence length settings, and its improvement over baseline models was even more significant in long-sequence scenarios.
Model Complexity and Runtime Analyses. To evaluate the overhead of WEARec, we assessed the number of parameters and runtime per epoch during training at $N=200$. The results are presented in Table 2. We can observe that WEARec exhibits a shorter training time compared to baseline models with competitive performance. Overall, WEARec’s total parameters are increased due to the use of a simple MLP. However, by not employing contrastive learning and self-attention mechanisms, WEARec actually trains faster than SLIME4Rec and BSARec. For more descriptions, interested readers should refer to Appendix B.4.
@@
(a) $k$
(b) $\alpha$
Figure 5: Performance of WEARec on HR@20 with varying hyperparameters.
Sensitivity to $k$. Figure 5(a) shows the HR@20 by varying $k$. The results indicate that an optimal $k$ value, neither too large nor too small, is critical for learning user interest preferences and consequently improving model performance.
Sensitivity to $\alpha$. Figure 5(b) shows the HR@20 by varying $\alpha$. These results suggest that optimal performance is more probable when $\alpha$ is approximately 0.3.
@@
Figure 6: Visualization of spectral responses for different types of filter models across layers in Beauty. More in-depth model analysis in Appendix C.
@@
Early SR research often relied on Markov chain assumptions (Rendle, Freudenthaler, and Schmidt-Thieme 2010). With the widespread adoption of deep learning methods (He et al. 2017), numerous studies have employed neural network architectures as encoders. Caser (Tang and Wang 2018) utilizes convolutional operations to capture higher-order patterns. SASRec (Kang and McAuley 2018) leverages self-attention mechanisms to capture item-item relationships. Furthermore, recent studies have enhanced sequential embedding representations through contrastive learning ($e.g.$, CL4SRec (Xie et al. 2022) and DuoRec (Qiu et al. 2022)). However, these time-domain models still struggle to effectively capture users’ underlying periodic behavioral patterns.
@@
Recently, researchers have begun applying frequency-domain analysis to sequential recommendation. FMLPRec (Zhou et al. 2022) pioneered frequency-based MLP filtering to capture periodic patterns. SLIME4Rec (Du et al. 2023a) and FEARec (Du et al. 2023b) further advanced this direction by proposing a layered frequency ramp structure integrated with contrastive learning. BSARec (Shin et al. 2024) sought to uncover fine-grained sequential patterns and inject them as inductive biases into the model. FamouSRec (Zhang et al. 2025) developed a MoE architecture for selecting specialized expert models tailored to users’ specific frequency-based behavioral patterns. However, these models either lack user-specific adaptivity or incur substantial computational costs.
@@ B.1 Details of Datasets
  • MovieLens-1M (Harper and Konstan 2015) is based on reviews of movies that were collected from the non-commercial movie recommendation website MovieLens. The interaction number in ML-1M is about 1 million.
  • Amazon Beauty, and Sports (McAuley et al. 2015) These datasets contain user-item interactions from the Amazon review dataset. They consist of product reviews and ratings, providing a rich data source for evaluating sequential recommendation models in the e-commerce domain.
@@ B.2 Details of Baselines
  • Time-domain SR models: GRU4Rec (Hidasi et al. 2015) is the first model to apply Gated Recurrent Unit (GRU) to model sequences of user behavior for sequential recommendation, Caser (Tang and Wang 2018) is a CNN-based method capturing local dynamic patterns of user activity by using horizontal and vertical convolutional filters over time, SASRec (Kang and McAuley 2018) captures relations between items in a sequence using the self-attention mechanism, and DuoRec (Qiu et al. 2022) uses unsupervised model-level augmentation and supervised semantic positive samples for contrastive learning.
  • Frequency-domain SR models: FMLPRec (Zhou et al. 2022) is a all-MLP model using a learnable filter enchanced block to remove noise in the embedding matrix, FamouSRec (Zhang et al. 2025) uses a Mixture-of-Experts (MoE) approach, allowing the model to focus on different frequency ranges via heterogeneous encoder modules, FEARec (Du et al. 2023b) utilizes frequency domain information in atten- tion computation and integrates information from both time and frequency domains, SLIME4Rec (Du et al. 2023a) utilizes a frequency ramp structure to consider different frequency bands for each layer with dynamic and static selection modules, BSARec (Shin et al. 2024) adjusts the influence on the high-frequency region to be learnable and utilizes it as an inductive bias of self-attention. It is the most recent and strong baseline for sequential recommendation.

### v2

@@ Abstract
Sequential recommendation has garnered significant attention for its ability to capture dynamic preferences by mining users’ historical interaction data. Given that users’ complex and intertwined periodic preferences are difficult to disentangle in the time domain, recent research is exploring frequency domain analysis to identify these hidden patterns. However, current frequency-domain-based methods suffer from two key limitations: (i) They primarily employ static filters with fixed characteristics, overlooking the personalized nature of behavioral patterns; (ii) While the global discrete Fourier transform excels at modeling long-range dependencies, it can blur non-stationary signals and short-term fluctuations. To overcome these limitations, we propose a novel method called Wavelet Enhanced Adaptive Frequency Filter for Sequential Recommendation (WEARec). Specifically, it consists of two vital modules: dynamic frequency-domain filtering and wavelet feature enhancement. The former is used to dynamically adjust filtering operations based on behavioral sequences to extract personalized global information, and the latter integrates wavelet transform to reconstruct sequences, enhancing blurred non-stationary signals and short-term fluctuations. Finally, these two modules work synergistically to achieve comprehensive performance and efficiency optimization in long sequential recommendation scenarios. Extensive experiments on four widely-used benchmark datasets demonstrate the superiority of WEARec.
Codehttps://github.com/xhy963319431/WEARec
@@ Introduction
Sequential Recommendation (SR) plays a crucial role in e-commerce applications by capturing users’ dynamic interest shifts through their historical interaction data (music; music2). The remarkable success of the transformer architecture in Natural Language Processing (NLP) (attention) and Computer Vision (CV) (vit) has led to significant advancements in sequential recommendation (sasrec; feature; contrastive). This has directly inspired a multitude of sequential recommendation models based on self-attention (bert; duo; s3). However, items in user interactions are typically chronologically entangled and inherently noisy (slime; fea). Consequently, it is challenging for models to directly discern changes in behavioral preferences from raw sequences within the temporal domain (fmlp).
@@
To address this limitation, recent research has begun exploring frequency-domain approaches to replace self-attention mechanisms (gfnet; fft; glfnet). By decomposing user sequences into different frequency components ($e.g.$, high-frequency and low-frequency signals) using Fourier transform, periodic patterns that are difficult to identify in the time domain can be effectively captured (fmlp; slime; bsa; famous). For example, FMLPRec (fmlp) pioneered frequency domain processing of sequential data, replacing the self-attention mechanism with learnable filters. SLIME4Rec (slime) proposed a frequency ramp structure, which considers different frequency bands for each layer. BSARec (bsa) used a frequency domain re-tuning component as an inductive bias for self-attention.
@@
However, despite their success in SR, existing frequency-domain sequential recommendation models have two key limitations. First, existing methods typically apply a static, fixed-pattern filter to all frequency components, which uniformly processes all user sequences, ignoring the personalized nature of behavioral patterns (bsa; famous). This unified filtering approach is susceptible to the influence of dominant users in the dataset and fails to consider the characteristics of individual users. In fact, users often exhibit diverse behavioral patterns: some user behaviors follow long-term preferences ($e.g.$, low-frequency signals), while other user behaviors show the opposite trend (slime; famous). To illustrate this, we trained a classic sequential recommendation model ($i.e.$, FMLPRec (fmlp)) on the Beauty and Sports datasets, and replaced its learnable filters with specific band-pass filters. By statistically analyzing how many users could be correctly predicted by specific frequency components, we could identify which users were driven by particular frequency components. The results, shown in Figure 1, indicate that users exhibit diverse behavioral patterns, with each focusing on different frequencies. This emphasizes the importance of developing personalized filtering models to capture individualized user behavioral patterns.
@@
The second limitation is related to the low-pass filtering characteristic of frequency-domain filters. The Discrete Fourier Transform (DFT) analyzes signal components globally, primarily serving as a global, rather than local, frequency extraction method. While the global DFT excels at capturing long-range dependencies in current frequency-domain recommendation models (fmlp; slime), it struggles to capture the local temporal features of high-frequency interactions and short-term points of interest (wave). For instance, FMLPRec has been shown to essentially act as a low-pass filter (bsa). Although SLIME4Rec attempts to balance high-frequency and low-frequency representations through its hierarchical learning mechanism, the model still tends to learn low-frequency components within the hierarchical frequency bands.
@@
To overcome these challenges, we propose Wavelet Enhanced Adaptive Frequency Filter for Sequential Recom- mendation (WEARec). WEARec consists of two key modules: Dynamic Frequency Filtering (DFF) and Wavelet Feature Enhancement (WFE) module. Specifically, the DFF module uses a simple Multi-Layer Perceptron (MLP) to enhance or suppress specific frequency bands based on context signals, ensuring effective global fusion. Furthermore, the WFE module reconstructs sequences via wavelet transform, amplifying obscure non-stationary signals and short-term fluctuations that are prone to being blurred by global DFT. Finally, to ensure all frequency components are considered and to best preserve meaningful periodic user features, we blend the DFF module with the WFE module. Moreover, our proposed model achieves better performance with lower computational costs, especially in long-sequence scenarios. The main contributions of this paper are summarized as follows:
@@
Where $\tilde{L}$ and $\tilde{H}$ are reconstruction filters. In our paper, the forward DWT converts sequential behavior into high/low-frequency sub-signals and denote it by $\mathcal{W}(\textperiodcentered)$. The IDWT (denoted by $\mathcal{W}^{-1}(\textperiodcentered)$) reconstructs decomposed sub-signals into the original signal. For more descriptions, interested readers should refer to Appendix A (arxiv).
@@
This module captures fine-grained temporal patterns through differentiable wavelet transforms. Here, the Haar wavelet transform (haar) was selected due to its simple structure, high computational efficiency, and the desirable property of perfect signal reconstruction.
@@
In the final layer of WEARec, we can compute the recommendation probability for each candidate item to predict how likely the user would adopt the item. Specifically, the corresponding predicted probability $\hat{\mathbb{y}}$ can be generated by:
@@
where $\hat{\mathbb{y}}\in\mathbb{R}^{|\mathcal{V}|}$. and $\mathbb{h}^{L}\in\mathbb{R}^{1\times d}$ is the output of the $L$-layer blocks at the final step. To optimize the model parameters, we therefore use the cross-entropy loss (duo; fea; slime; bsa). The objective function of SR can be formulated as:
@@
Datasets. We conduct experiments on four public datasets collected from real-world platforms in order to thoroughly evaluate WEARec. i,ii) Beauty and Sports from Amazon (amazon), iii) ML-1M (movielens), iv) LastFM. Following (s3; fmlp) , we also adopt the 5-core settings by filtering out users with less than 5 interactions. The detailed dataset statistics are presented in Appendix B.1 (arxiv).
Evaluation Metrics. In our evaluation, we adopt the leave-one-out strategy for partitioning each user’s item sequence (s3). We rank the prediction scores throughout the entire item set without using negative sampling, as recommended by (sampled). Performance is evaluated on a variety of evaluation metrics, including Hit Ratio at $K$ (HR@$K$) and Normalized Discounted Cumulative Gain at $K$ (NDCG@$K$, NG@$K$) on all datasets. The $K$ is set to 10 and 20.
@@
Time-domain SR models: GRU4Rec (gru), Caser (caser), SASRec (sasrec), and DuoRec (duo).
Frequency-domain SR models: FMLPRec (fmlp), FamouSRec (famous), FEARec (fea), SLIME4Rec (slime), BSARec (bsa).
Implementation Details. We implement our WEARec model in PyTorch. For the baseline models, we refer to their best hyper-parameters setups reported in the original papers and directly report their reimplementations results if available, since the datasets and evaluation metrics used in these works are strictly consistent with ours. Both the dimension of the feed-forward network and item embedding size are set to 64. The number of WEARec blocks $L$ is set to 2, and the maximum sequence length $N$ is set to 50. Batch size is set to 256. The model is optimized by Adam optimizer with a learning rate from {0.0005,0.001}. The wavelet decomposition level is set to 1. The $\alpha$ is in {0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9}, and the $k$ chosen from {1,2,4,8}. We report the result of each model under its optimal hyper-parameter settings. The best hyperparameters are in Appendix B.3 (arxiv) for reproducibility.
@@
Table 2: The number of parameters and training time (runtime per epoch) for $N=200$ on ML-1M and LastFM. More results are in Appendix B (arxiv).
Model performance. To investigate the impact of long sequence scenarios on recommendation results, we varied the maximum sequence length $N$ for FMLPRec, BSARec, SLIME4Rec and WEARec. We selected the LastFM and ML-1M datasets, which have longer average sequence lengths, for our experiments. Figure 3 presents the experimental results in terms of HR@20. We have obtained similar experimental results in terms of other metrics. We observed that almost all models achieved their best performance at $N=200$, indicating that longer sequence information can more comprehensively represent user behavior patterns. Furthermore, while baseline models showed performance improvements in long-sequence scenarios, they were prone to overfitting, leading to performance convergence. Finally, WEARec consistently outperformed the baselines across all different maximum sequence length settings, and its improvement over baseline models was even more significant in long-sequence scenarios. For more descriptions, interested readers should refer to Appendix B.4 (arxiv).
Model Complexity and Runtime Analyses. To evaluate the overhead of WEARec, we assessed the number of parameters and runtime per epoch during training at $N=200$. The results are presented in Table 2. We can observe that WEARec exhibits a shorter training time compared to baseline models with competitive performance. Overall, WEARec’s total parameters are increased due to the use of a simple MLP. However, by not employing contrastive learning and self-attention mechanisms, WEARec actually trains faster than SLIME4Rec and BSARec. For more descriptions, interested readers should refer to Appendix B.5 (arxiv).
@@
Figure 5: Performance of WEARec on HR@20 with varying hyperparameters..
Sensitivity to $k$. Figure 5 shows the HR@20 by varying $k$. The results indicate that an optimal $k$ value, neither too large nor too small, is critical for learning user interest preferences and consequently improving model performance.
Sensitivity to $\alpha$. Figure 5 shows the HR@20 by varying $\alpha$. These results suggest that optimal performance is more probable when $\alpha$ is approximately 0.3.
@@
Figure 6: Visualization of spectral responses for different types of filter models across layers in Beauty. More in-depth model analysis in Appendix C (arxiv).
@@
Early SR research often relied on Markov chain assumptions (fmc). With the widespread adoption of deep learning methods (ncf), numerous studies have employed neural network architectures as encoders. Caser (caser) utilizes convolutional operations to capture higher-order patterns. SASRec (sasrec) leverages self-attention mechanisms to capture item-item relationships. Furthermore, recent studies have enhanced sequential embedding representations through contrastive learning ($e.g.$, CL4SRec (cl4rec) and DuoRec (duo)). However, these time-domain models still struggle to effectively capture users’ underlying periodic behavioral patterns.
@@
Recently, researchers have begun applying frequency-domain analysis to sequential recommendation. FMLPRec (fmlp) pioneered frequency-based MLP filtering to capture periodic patterns. SLIME4Rec (slime) and FEARec (fea) further advanced this direction by proposing a layered frequency ramp structure integrated with contrastive learning. BSARec (bsa) sought to uncover fine-grained sequential patterns and inject them as inductive biases into the model. FamouSRec (famous) developed a MoE architecture for selecting specialized expert models tailored to users’ specific frequency-based behavioral patterns. However, these models either lack user-specific adaptivity or incur substantial computational costs.
@@ B.1 Details of Datasets
  • MovieLens-1M (movielens) is based on reviews of movies that were collected from the non-commercial movie recommendation website MovieLens. The interaction number in ML-1M is about 1 million.
  • Amazon Beauty, and Sports (amazon) These datasets contain user-item interactions from the Amazon review dataset. They consist of product reviews and ratings, providing a rich data source for evaluating sequential recommendation models in the e-commerce domain.
@@ B.2 Details of Baselines
  • Time-domain SR models: GRU4Rec (gru) is the first model to apply Gated Recurrent Unit (GRU) to model sequences of user behavior for sequential recommendation, Caser (caser) is a CNN-based method capturing local dynamic patterns of user activity by using horizontal and vertical convolutional filters over time, SASRec (sasrec) captures relations between items in a sequence using the self-attention mechanism, and DuoRec (duo) uses unsupervised model-level augmentation and supervised semantic positive samples for contrastive learning.
  • Frequency-domain SR models: FMLPRec (fmlp) is a all-MLP model using a learnable filter enchanced block to remove noise in the embedding matrix, FamouSRec (famous) uses a Mixture-of-Experts (MoE) approach, allowing the model to focus on different frequency ranges via heterogeneous encoder modules, FEARec (fea) utilizes frequency domain information in atten- tion computation and integrates information from both time and frequency domains, SLIME4Rec (slime) utilizes a frequency ramp structure to consider different frequency bands for each layer with dynamic and static selection modules, BSARec (bsa) adjusts the influence on the high-frequency region to be learnable and utilizes it as an inductive bias of self-attention. It is the most recent and strong baseline for sequential recommendation.